OpenStudy (homeworksucks):

sin((pi)/2)+(theta)=cos(theta)

OpenStudy (anonymous):

do you mean this? \[\sin (\pi/2) + \theta = \cos (\theta)\]

OpenStudy (homeworksucks):

yeah

OpenStudy (homeworksucks):

without brackets i guess

OpenStudy (anonymous):

I think you may have copied it wrong because that's kind of impossible to solve...

OpenStudy (homeworksucks):

pi/2 as in radians

OpenStudy (homeworksucks):

I have to establish an identity s

OpenStudy (anonymous):

u're equation should be \[\sin (\frac{ \pi }{ 2 }+\theta) = \cos \theta\] otherwise there will be no meaning

OpenStudy (homeworksucks):

typo

OpenStudy (homeworksucks):

yeah

OpenStudy (anonymous):

so what is the question?

OpenStudy (homeworksucks):

your last reply

OpenStudy (anonymous):

okay do we have to prove or solve for theta

OpenStudy (homeworksucks):

prove

OpenStudy (anonymous):

okay to prove this kinds of equation we only develop the complex side until it is equal to the simple one okay

OpenStudy (anonymous):

do u knw double angles identity?

OpenStudy (homeworksucks):

yes

OpenStudy (anonymous):

okay use it to expand the left side of ur equation and see what u get

OpenStudy (homeworksucks):

2sin(a)+2cos(a)?

OpenStudy (homeworksucks):

thats the double angle formula right?

OpenStudy (homeworksucks):

sin(pi/2+theta)= cos theta

OpenStudy (homeworksucks):

sorry sin(a+b)=sinacosb+cosa+sinb?

OpenStudy (homeworksucks):

is that the one i want?

OpenStudy (anonymous):

s_sorry i used a wrong spelling it actually two angle. sin(a+b)=sinacosb+cosasinb use it to expand ur left hand side of ur equation

OpenStudy (anonymous):

yes but there is no plus sign between cosa and sinb

OpenStudy (homeworksucks):

so sin((pi/2)+theta)=sin pi/2+cos theta+cos pi/2 sin theta=1*cos theta+0*sin theta= cos theta?

OpenStudy (anonymous):

yes but just check ur plus signs again